OpenStudy (anonymous):

dz/dx when x=0,where z=In(cx+d) (c and d are constants,with d bigger than 0

OpenStudy (cristiann):

z is a composed function, so: dz/dx(x)=(cx+d)'/(cx+d)=c/(cx+d) dz/dx(0)=c/d

OpenStudy (anonymous):

thax loads